The value of the multiplier associated with changes in autonomous spending is given by k.
a) Find the value of the multiplier when MPC=0.50, 0.75 and 0.80.
b) What is the relationship between the value of the multiplier and MPC?
c) Find the change in equilibrium level of output when there is a k10 000 increase in investment spending and MPC=0.50, 0.75 AND 0.80.
1
Expert's answer
2020-06-16T14:01:37-0400

a) Multiplier formula;


"K=1\/(1-MPC)"


K is multiplier

MPC is marginal propensity to consume


"K=1\/(1-0.5)"

"K=2"


"K=1\/(1-0.75)"

"K=4"


"K=1\/(1-0.85)"

"K=6.7"


b)

Multiplier is directly related to marginal propensity to consume,the higher the multiplier,the higher the marginal propensity to consume and vice versa.


c)


Since the initial increase in spending is k 10,000 and the multiplier is 2, with Mpc of 0.5

the change in equilibrium output will be:

"2*10,000=20,000"


Spending is k 10,000 and the multiplier is 4, with Mpc of 0.75

"4*10,000=40,000"


Spending is k 10,000 and the multiplier is 6.7, with Mpc of 0.85

"6.7*10,000=67,000"